# GET: Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v1

@@ Abstract
Given unlabelled datasets containing both old and new categories, generalized category discovery (GCD) aims to accurately discover new classes while correctly classifying old classes, leveraging the class concepts learned from labeled samples. Current GCD methods only use a single visual modality of information, resulting in poor classification of visually similar classes. Though certain classes are visually confused, their text information might be distinct, motivating us to introduce text information into the GCD task. However, the lack of class names for unlabelled data makes it impractical to utilize text information. To tackle this challenging problem, in this paper, we propose a Text Embedding Synthesizer (TES) to generate pseudo text embeddings for unlabelled samples. Specifically, our TES leverages the property that CLIP can generate aligned vision-language features, converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings. Besides, we employ a dual-branch framework, through the joint learning and instance consistency of different modality branches, visual and semantic information mutually enhance each other, promoting the interaction and fusion of visual and text embedding space. Our method unlocks the multi-modal potentials of CLIP and outperforms the baseline methods by a large margin on all GCD benchmarks, achieving new state-of-the-art. The code will be released at https://github.com/enguangW/GET.
Keywords:
Generalized Category Discovery
@@ 1 Introduction
Previous GCD methods [36, 40, 43, 30, 45] utilize a DINO [4] pre-trained ViT as the backbone network to expect good initial discrimination ability of the model, thereby facilitating fine-tuning on the training data. Although promising results have been achieved, the representation of the DINO pre-trained model focuses on specific aspects of the data, resulting in large differences in the clustering performance for different classes in the downstream dataset. To explore the impact of different types of pre-trained models on GCD clustering performance, we conduct an extensive empirical investigation in Sec. 5.4. A key observation is that various pre-trained models perform inferior on distinguishing certain classes, especially those with visually similar appearances, such as the classes in all fine-grained datasets and some super-class subsets of generic datasets. We argue that this is due to the reliance on single visual modality information. Current GCD methods only utilize a single visual modality of information, even visual features obtained by the powerful CLIP [32] visual backbone are still difficult to generalize certain visual concepts, leading to sub-optimal results (see in Fig. 1 (a)). An evident fact is that, even if certain classes are visually confused, their textual information might be different and more discriminative, motivating us to introduce text information into the GCD task.
@@
As we all know, thanks to the cross-modal alignment training of very large-scale image-text pairs, CLIP shows strong zero-shot performance, demonstrates its powerful generalization ability of multi-modal joint embedding, and a wide range of fine-grained visual concepts can be well separated in the multi-modal embedding space. In order to utilize multi-modal joint embedding for GCD tasks, it is a natural idea to conduct joint clustering training with the visual embedding of images and the text embedding of the corresponding class names. However, the lack of class names for unlabelled data in GCD makes it impractical to use the text encoder, thus locking the multi-modal potential of CLIP.
@@
To tackle this challenging problem, in this paper, we propose a generative-based method to GEnerate pseudo Text embeddings for unlabelled data, dubbed GET. In particular, we introduce a Text Embedding Synthesizer (TES) module based on the vision-language alignment property of CLIP, producing reliable and modality-aligned pseudo text features. As shown in Fig. 1 (b), TES learns a mapping that transforms image embeddings into text embeddings. Specifically, the TES module converts visual embeddings into tokens of the CLIP’s text encoder, eliminating the need for textual input. To ensure that generated pseudo text embeddings reside in the same text space as real text embeddings and maintain consistency, TES aligns them with the real text embeddings generated by corresponding class names for labeled data, thus leading to an implicit alignment for unlabelled data. Besides, TES simultaneously aligns the text embedding and the image embedding of the same instance, enforcing the consistency between language and vision while preventing overfitting to known classes. This training approach renders TES equivalent to a special finetuned CLIP text encoder on a specific dataset with only visual input. From another perspective, our TES can be considered as performing an image captioning task [28].
@@
To summarize, our contributions are as follows: (i) Through extensive empirical research, we demonstrate that GCD methods are constrained by pre-trained backbone and the representation of a single visual modality makes current GCD methods have inferior clustering performance on many visually similar classes. (ii) To unlock the potential of CLIP, We propose a TES module converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings for each sample. Besides, through the cross-modal instance consistency objective in our dual-branch framework, information of different modalities mutually enhances each other, producing more discriminative classification prototypes. (iii)Our method achieves state-of-the-art results on multiple benchmarks, extending GCD to a multi-modal paradigm.
@@ 2 Related Works
Novel Class Discovery.
@@
Generalized Category Discovery.
@@
Vision-Language Pre-training.
@@ 3.1 Problem formulation
In the context of GCD, the training data $\mathcal{D}$ is divided into a labeled dataset $\mathcal{D}_{l}=\left\{(\boldsymbol{x}_{i}^{l},{y}_{i}^{l})\right\}_{i=1}^{N}\in\mathcal{X}\times\mathcal{Y}_{l}$ and an unlabelled dataset $\mathcal{D}_{u}=\left\{(\boldsymbol{x}_{i}^{u},{y}_{i}^{u})\right\}_{i=1}^{M}\in\mathcal{X}\times\mathcal{Y}_{u}$, where $\mathcal{Y}_{l}$ and $\mathcal{Y}_{u}$ represent the label space while $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$, and $\mathcal{D}=\mathcal{D}_{l}\cup\mathcal{D}_{u}$. $|\mathcal{Y}_{l}|$ and $|\mathcal{Y}_{u}|$ represent the number of categories for labeled samples and unlabelled samples, respectively. The goal of GCD is to correctly cluster unlabelled samples with the help of labeled samples.
@@ 3.2 Parametric GCD method (SimGCD)
In this paper, we tackle the GCD problem in a parametric way which is proposed by SimGCD [40]. It trains a unified prototypical classification head for all new/old classes to perform GCD through a DINO-like form of self-distillation. Specifically, it includes two types of loss functions: representation learning and parametric classification. For representation learning, it performs supervised representation learning [18] on all labeled data and self-supervised contrastive learning on all training data, the loss functions are as follows:
@@
where $\mathcal{N}_{i}$ denotes the indices of other images with the same semantic label as $\boldsymbol{x}_{i}$ in a batch, $B_{l}$ corresponds to the labeled subset of the mini-batch $B$ and $\tau_{c}$ is a temperature value. For visual embeddings $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ of two views $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i}^{\prime}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ to high-dimensional embeddings $\boldsymbol{h}_{i}=g(\boldsymbol{z}_{i})$ and $\boldsymbol{h}_{i}^{\prime}=g(\boldsymbol{z}_{i}^{\prime})$. For parametric classification, all labeled data are trained by a cross-entropy loss and all training data are trained by a self-distillation loss:
@@ 4 Our Method
In this paper, we propose GET, which addresses the GCD task in a multi-modal paradigm. As shown in Fig. 2, our GET contains two stages. In the first stage, we learn a text embedding synthesizer (TES, in Sec. 4.1) to generate pseudo text embeddings for each sample. In the second stage, a dual-branch multi-modal joint training strategy with cross-modal instance consistency (in Sec. 4.2) is introduced to fully leverage multi-modal features.
@@ 4.1 Text embedding synthesizer
The absence of natural language class names for unlabeled data makes it challenging to introduce language information into the GCD task. In this paper, we attempt to generate pseudo text embeddings aligned with visual embeddings for each image from a feature-based perspective.
Inspired by BARON [41], which treats embeddings within bounding boxes as embeddings of words in a sentence to solve the open-vocabulary detection task, we propose a text embedding synthesizer (TES). Specifically, our TES leverages the property that CLIP can generate aligned vision-language features, converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings for each sample. The architecture of TES is shown in Fig. 2 (a). For each image $\boldsymbol{x}_{i}$ in a mini-batch, we use CLIP’s image encoder to obtain its visual embedding $\boldsymbol{z}_{i}^{v}$. A single fully connected layer is used to map the visual embedding to pseudo tokens that serve as input to the CLIP’s text encoder, thus generating corresponding pseudo text embeddings $\boldsymbol{\hat{z}}_{i}^{t}$.
@@
$\mathcal{L}_{align}^{v}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left({\boldsymbol{z}_{i}^{v}}^{\top}{\boldsymbol{\hat{z}}_{i}^{t}}/\tau_{a}\right)}{\sum_{j\in B}\exp\left({\boldsymbol{z}_{i}^{v}}{\boldsymbol{\hat{z}}_{j}^{t}}/\tau_{a}\right)}\,,$ (5)
$\mathcal{L}_{align}^{t}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left({\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{\boldsymbol{z}_{i}^{v}}/\tau_{a}\right)}{\sum_{j\in B}\exp\left({\boldsymbol{\hat{z}}_{i}^{t}},{\boldsymbol{z}_{j}^{v}}/\tau_{a}\right)}\,,$ (6)
@@
$\mathcal{L}_{distill}=-\frac{1}{|B_{l}|}\sum_{i\in B_{l}}\left(\log\frac{\exp\left({{\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{T(n_{i})}}\right)}{\sum_{j=0}^{|\mathcal{Y}_{l}|}\mathbbm{1}_{[j\neq n_{i}]}\exp\left({\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{T(j)}\right)}+\left(\boldsymbol{\hat{z}}_{i}^{t}-T(n_{i})\right)^{2}\right)\,,$ (7)
@@
By aligning the pseudo text embeddings of labeled data and their corresponding real text embeddings, the pseudo text embeddings for unlabelled data can also implicitly align with the real embeddings. The overall objective of our text embedding synthesizer is calculated as:
@@
Moreover, we introduce a multi-view strategy for TES. Specifically, we calculate both the align loss $\mathcal{L}_{align}$ and distill loss $\mathcal{L}_{distill}$ for two different views $\boldsymbol{x}_{i}$ and ${\boldsymbol{x}_{i}}^{\prime}$ of the same image in a mini-batch. This further implicitly enhances the instance discriminative nature [42] of our TES training. This allows different views of the same image to generate identical pseudo text embeddings while aligning with their shared visual features. The generated pseudo text embeddings $\boldsymbol{\hat{z}}_{i}^{t}$ are then used for joint training in the second stage.
@@ 4.2 Dual-branch multi-modal joint training
Intuitively, the introduction of multi-modal information can have a positive impact on the GCD task. Categories with similar visual concepts may be distinguished on the basis of textual concepts. However, how to effectively utilize visual and text information in the GCD task and make the most of their respective roles remains challenging. In this paper, we propose a dual-branch architecture as illustrated in Fig. 2 (b), which focuses on semantic and visual information, respectively. We employ the same parametric training strategy (in Sec. 3.2) for each branch to promote that the model has aligned and complementary discriminative capabilities for visual and text features of the same class. Furthermore, we introduce a cross-modal instance consistency loss, which constrains the instance relationships of samples in both visual and text spaces, enabling the two branches to learn from each other. We use the indicator $v$ to represent the visual concept while $t$ for the text concept.
Visual-branch.
The objective of the visual branch contains a representation learning part and a parametric classification part. Given a visual embedding $\boldsymbol{z}_{i}^{v}$ of image $\boldsymbol{x}_{i}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}^{v}$ to a high-dimensional embedding $\boldsymbol{h}_{i}^{v}=g(\boldsymbol{z}_{i}^{v})$. Meanwhile, we employ a prototypical classifier $\eta(\cdot)$ to generate classification probability distribution $\boldsymbol{p}_{i}^{v}=\eta(\boldsymbol{z}_{i}^{v})$. Simply replace all the high-dimensional embedding $\boldsymbol{h}$ in Eqs. 1 and 2 with its corresponding visual branch version $\boldsymbol{h}^{v}$ can obtain the the supervised contrastive loss $\mathcal{L}_{scon}^{v}$ and the self-supervised contrastive loss $\mathcal{L}_{ucon}^{v}$. The overall representation learning loss is balanced with $\lambda$, written as:
@@
For the parametric classification part, just replace $\boldsymbol{p}_{i}$ and ${\boldsymbol{p}_{i}}^{\prime}$ of Eq. 3 and Eq. 4 with $\boldsymbol{p}_{i}^{v}$ and ${\boldsymbol{p}_{i}^{v}}^{\prime}$ can obtain the cross-entropy loss $\mathcal{L}_{\text{cls-v}}^{s}$ and the self-distillation loss $\mathcal{L}_{\text{cls-v}}^{u}$. Thus, the classification loss is $\mathcal{L}_{\text{cls}}^{v}=(1-\lambda)\mathcal{L}_{\text{cls-v}}^{u}+\lambda\mathcal{L}_{\text{cls-v}}^{s}$.
@@
$\mathcal{L}_{\text{dual}}^{v}=\mathcal{L}_{\text{rep}}^{v}+\mathcal{L}_{\text{cls}}^{v}\,.$ (10)
Text-branch.
Our text branch simply adopts the same training strategy as the visual branch. That is, in particular, given a text embedding $\boldsymbol{\hat{z}}_{i}^{t}$ generated by TES, we first input it into a fully connected layer to gain a learnable text embedding $\boldsymbol{\hat{z}}_{i}^{tl}$ while change its dimension. Simply replace $\boldsymbol{h}_{i}^{v}$ in the representation learning objective $\mathcal{L}_{\text{rep}}^{v}$ with $\boldsymbol{h}_{i}^{t}=g(\boldsymbol{\hat{z}}_{i}^{tl})$ and replace ${\boldsymbol{p}_{i}^{v}}$ in classification parts $\mathcal{L}_{\text{cls}}^{v}$ with $\boldsymbol{p}_{i}^{t}=\eta(\boldsymbol{\hat{z}}_{i}^{tl})$ can yield the corresponding text objectives $\mathcal{L}_{\text{rep}}^{t}$ and $\mathcal{L}_{\text{cls}}^{t}$. In other words, changing the visual conception indicator $v$ into text conception indicator $t$ can get the corresponding objectives for the text branch. Thus, the overall objective of our text branch can be formally written as:
$\mathcal{L}_{\text{dual}}^{t}=\mathcal{L}_{\text{rep}}^{t}+\mathcal{L}_{\text{cls}}^{t}\,.$ (11)
@@
Cross-modal instance consistency objective.
In order to enable the two branches to learn from each other while encouraging agreement between two different modes, we propose a cross-modal instance consistency objective (CICO), shown in Fig. 2 (c). Our CICO has the same form of mutual knowledge distillation as  [44], but we distill the instance consistency between the two branches. For each mini-batch $B$, we choose its labeled subset $B_{l}$ containing $K$ categories as anchor samples, calculate the visual and text prototypes for $K$ categories as visual anchors $\mathcal{P}_{v}$ and text anchors $\mathcal{P}_{t}$, respectively. We define the instance relationships in visual and text branch as:
$s_{i}^{v}=\sigma(sim(\boldsymbol{z}_{i}^{v},\mathcal{P}_{v})\,,$ (12)
$s_{i}^{t}=\sigma(sim(\boldsymbol{\hat{z}}_{i}^{tl},\mathcal{P}_{t}))\,.$ (13)
@@
where $D_{KL}$ is the Kullback-Leibler divergence. Mutual knowledge distillation on instance relationships for two modalities makes visual and text flows exchange and benefit from each other, which is demonstrated to be effective in Sec. 5.3.

### v2

# Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

@@ Abstract
Given unlabelled datasets containing both old and new categories, generalized category discovery (GCD) aims to accurately discover new classes while correctly classifying old classes, leveraging the class concepts learned from labeled samples. Current GCD methods only use a single visual modality of information, resulting in poor classification of visually similar classes. As a different modality, text information can provide complementary discriminative information, which motivates us to introduce it into the GCD task. However, the lack of class names for unlabelled data makes it impractical to utilize text information. To tackle this challenging problem, in this paper, we propose a Text Embedding Synthesizer (TES) to generate pseudo text embeddings for unlabelled samples. Specifically, our TES leverages the property that CLIP can generate aligned vision-language features, converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings. Besides, we employ a dual-branch framework, through the joint learning and instance consistency of different modality branches, visual and semantic information mutually enhance each other, promoting the interaction and fusion of visual and text knowledge. Our method unlocks the multi-modal potentials of CLIP and outperforms the baseline methods by a large margin on all GCD benchmarks, achieving new state-of-the-art. The code will be released at https://github.com/enguangW/GET.
Index Terms:
Generalized category discovery, clustering, classification, vision-language pre-training
@@ 1 Introduction
Previous GCD methods [3, 4, 5, 6, 7] utilize a DINO [8] pre-trained ViT as the backbone network to expect good initial discrimination ability of the model, thereby facilitating fine-tuning on the training data. Although promising results have been achieved, the representation of the DINO pre-trained model focuses on specific aspects of the data, resulting in large differences in the clustering performance for different classes in the downstream dataset. To explore the impact of different types of pre-trained models on GCD clustering performance, we conduct an extensive empirical investigation in §$\S$5.5. A key observation is that various pre-trained models perform inferior on distinguishing certain classes, especially those with visually similar appearances, such as the classes in all fine-grained datasets and some super-class subsets of generic datasets. We argue that this is due to the reliance on single visual modality information. Current GCD methods only utilize a single visual modality of information, even visual features obtained by the powerful CLIP [9] visual backbone are still difficult to generalize certain visual concepts, leading to sub-optimal results (see in Fig. 1 (a)). Inspired by the idea that the textual modality can provide complementary discriminative information, we decided to introduce text information into the GCD task to compensate for the insufficient discriminative of visual concepts.
@@
As we all know, thanks to the cross-modal alignment training of very large-scale image-text pairs, CLIP shows strong zero-shot performance, and demonstrates its powerful generalization ability of multi-modal joint embedding. In order to utilize multi-modal joint embedding for GCD tasks, it is a natural idea to conduct joint clustering training with the visual embedding of images and the text embedding of the corresponding class names. However, the lack of class names for unlabelled data in GCD makes it impractical to use the text encoder, thus locking the multi-modal potential of CLIP for GCD.
@@
To tackle this challenging problem, in this paper, we propose a generative-based method to GEnerate pseudo Text embeddings for unlabelled data, dubbed GET. In particular, we introduce a Text Embedding Synthesizer (TES) module based on the vision-language alignment property of CLIP, producing reliable and modality-aligned pseudo text features. As shown in Fig. 1 (b), TES learns a mapping that transforms image embeddings into text embeddings. Specifically, the TES module converts visual embeddings into tokens of the CLIP’s text encoder, eliminating the need for textual input. To ensure that generated pseudo text embeddings reside in the same text space as real text embeddings and maintain consistency, TES distill knowledge from the real text embeddings generated by corresponding class names for labeled data, thus leading to an implicit alignment for unlabelled data. Besides, TES simultaneously aligns the text embedding and the image embedding of the same instance, enforcing the consistency between language and vision while preventing overfitting to known classes. This training approach renders TES equivalent to a special finetuned CLIP text encoder on a specific dataset with only visual input. From another perspective, our TES can be considered as performing an image captioning task [10].
@@
To summarize, our contributions are as follows:
  • Through extensive empirical research, we demonstrate that GCD methods are constrained by pre-trained backbone and the representation of a single visual modality makes current GCD methods have inferior clustering performance on many visually similar classes.
  • To tackle the problem that the text encoder can not be used on the unlabelled data, we propose a TES module converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings for each sample. Besides, through the cross-modal instance consistency objective in our dual-branch framework, information of different modalities mutually enhances each other, producing more discriminative classification prototypes.
  • Our method achieves state-of-the-art results on multiple benchmarks, extending GCD to a multi-modal paradigm. Meanwhile, we evaluated GCD methods on ImageNet-R, demonstrating that GCD based on multi-modal features can generalize across different domains, challenging the traditional assumption that the GCD task is based on the same domain.
@@ 2 Related Works
In this section, we first introduce the task of novel class discovery, followed by a review of the existing research on generalized category discovery. Since our proposed method uses the vision-language pre-trained model, we also briefly survey the development of vision-language pre-training.
2.1 Novel Class Discovery.
@@
2.2 Generalized Category Discovery.
@@
2.3 Vision-Language Pre-training.
@@ 3.1 Problem formulation
In the context of GCD, the training data $\mathcal{D}$ is divided into a labeled dataset $\mathcal{D}_{l}=\left\{(\boldsymbol{x}_{i}^{l},{y}_{i}^{l})\right\}_{i=1}^{N}\in\mathcal{X}\times\mathcal{Y}_{l}$ and an unlabelled dataset $\mathcal{D}_{u}=\left\{(\boldsymbol{x}_{i}^{u},{y}_{i}^{u})\right\}_{i=1}^{M}\in\mathcal{X}\times\mathcal{Y}_{u}$, where $\mathcal{Y}_{l}$ and $\mathcal{Y}_{u}$ represent the label space while $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$, and $\mathcal{D}=\mathcal{D}_{l}\cup\mathcal{D}_{u}$. $|\mathcal{Y}_{l}|$ and $|\mathcal{Y}_{u}|$ represent the number of categories for labeled samples and unlabelled samples, respectively. Following the setting in [3, 4], we assume the class number of new classes $|\mathcal{Y}_{u}\backslash\mathcal{Y}_{l}|$ is known, or it can be estimated through some off-the-shelf methods [3, 2]. The goal of GCD is to correctly cluster unlabelled samples with the help of labeled samples.
@@ 3.2 Parametric GCD method (SimGCD)
In this paper, we tackle the GCD problem in a parametric way which is proposed by SimGCD [4]. It trains a unified prototypical classification head for all new/old classes to perform GCD through a DINO-like form of self-distillation. Specifically, it includes two types of loss functions: representation learning and parametric classification. For representation learning, it performs supervised representation learning [26] $\mathcal{L}_{scon}$ on all labeled data and self-supervised contrastive learning $\mathcal{L}_{con}$ on all training data, the loss functions are as follows:
@@
where $\mathcal{N}_{i}$ denotes the indices of other images with the same semantic label as $\boldsymbol{x}_{i}$ in a batch, $B_{l}$ corresponds to the labeled subset of the mini-batch $B$ and $\tau_{c}$ is a temperature value. For visual embeddings $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ of two views $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i}^{\prime}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ to high-dimensional embeddings $\boldsymbol{h}_{i}=g(\boldsymbol{z}_{i})$ and $\boldsymbol{h}_{i}^{\prime}=g(\boldsymbol{z}_{i}^{\prime})$. For parametric classification, all labeled data are trained by a cross-entropy loss $\mathcal{L}_{\text{cls}}^{s}$ and all training data are trained by a self-distillation loss $\mathcal{L}_{\text{cls}}^{u}$:
@@ 4 Our Method
In this paper, we propose GET, which addresses the GCD task in a multi-modal paradigm. As shown in Fig. 2, our GET contains two stages. In the first stage, we learn a text embedding synthesizer (TES, in §$\S$4.1) to generate pseudo text embeddings for each sample. In the second stage, a dual-branch multi-modal joint training strategy with cross-modal instance consistency (in §$\S$4.2) is introduced to fully leverage multi-modal features.
@@ 4.1 Text embedding synthesizer
The absence of natural language class names for unlabelled data makes it challenging to introduce text information into the GCD task. In this paper, we attempt to generate pseudo text embeddings aligned with visual embeddings for each image from a feature-based perspective.
Inspired by BARON [28], which treats embeddings within bounding boxes as embeddings of words in a sentence to solve the open-vocabulary detection task, we propose a text embedding synthesizer (TES). Specifically, our TES leverages the property that CLIP can generate aligned vision-language features, converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings for each sample. The architecture of TES is shown in Fig. 2 (a). For each image $\boldsymbol{x}_{i}$ in a mini-batch, we use CLIP’s image encoder to obtain its visual embedding $\boldsymbol{z}_{i}^{v}$. A single fully connected layer $l$ is used to map the visual embedding to pseudo tokens that serve as input to the CLIP’s text encoder, thus generating corresponding pseudo text embeddings $\boldsymbol{\hat{z}}_{i}^{t}$.
@@
$\mathcal{L}_{align}^{v}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left({\boldsymbol{z}_{i}^{v}}^{\top}{\boldsymbol{\hat{z}}_{i}^{t}}/\tau_{a}\right)}{\sum_{j\in B}\exp\left({\boldsymbol{z}_{i}^{v}}^{\top}{\boldsymbol{\hat{z}}_{j}^{t}}/\tau_{a}\right)}\,,$ (5)
$\mathcal{L}_{align}^{t}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left({\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{\boldsymbol{z}_{i}^{v}}/\tau_{a}\right)}{\sum_{j\in B}\exp\left({\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{\boldsymbol{z}_{j}^{v}}/\tau_{a}\right)}\,,$ (6)
@@
$\begin{split}\mathcal{L}_{distill}=&-\frac{1}{|B_{l}|}\sum_{i\in B_{l}}\log\frac{\exp\left({{\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{T(n_{i})}}\right)}{\sum_{j=0}^{|\mathcal{Y}_{l}|}\mathbbm{1}_{[j\neq n_{i}]}\exp\left({\boldsymbol{\hat{z}}_{i}^{t}}^{\top}{T(j)}\right)}\\ &+\frac{1}{|B_{l}|}\sum_{i\in B_{l}}\left(\boldsymbol{\hat{z}}_{i}^{t}-T(n_{i})\right)^{2}\,,\end{split}$ (7)
@@
The overall objective of our text embedding synthesizer is calculated as:
@@
Moreover, we introduce a multi-view strategy for TES. Specifically, we calculate both the align loss $\mathcal{L}_{align}$ and distill loss $\mathcal{L}_{distill}$ for two different views $\boldsymbol{x}_{i}$ and ${\boldsymbol{x}_{i}}^{\prime}$ of the same image in a mini-batch. This further implicitly enhances the instance discriminative nature [29] of our TES training, allowing different views of the same labeled image to generate identical pseudo text embeddings. The generated pseudo text embeddings $\boldsymbol{\hat{z}}_{i}^{t}$ are then used for joint training in the second stage.
@@ 4.2 Dual-branch multi-modal joint training
Intuitively, the introduction of multi-modal information can have a positive impact on the GCD task. Textual information can serve as an effective complement to visual information, enhancing the model’s discriminative capabilities. However, how to effectively utilize visual and text information in the GCD task and make the most of their respective roles remains challenging. In this paper, we propose a dual-branch architecture as illustrated in Fig. 2 (b), which focuses on semantic and visual information, respectively. We employ the same parametric training strategy (in §$\S$3.2) for each branch to promote that the model has aligned and complementary discriminative capabilities for visual and text features of the same class. Furthermore, we introduce a cross-modal instance consistency loss, which constrains the instance relationships of samples in both visual and text spaces, enabling the two branches to learn from each other. We use the indicator $v$ to represent the visual concept while $t$ for the text concept.
Visual-branch.  The objective of the visual branch contains a representation learning part and a parametric classification part. Given a visual embedding $\boldsymbol{z}_{i}^{v}$ of image $\boldsymbol{x}_{i}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}^{v}$ to a high-dimensional embedding $\boldsymbol{h}_{i}^{v}=g(\boldsymbol{z}_{i}^{v})$. Meanwhile, we employ a prototypical classifier $\eta(\cdot)$ to generate classification probability distribution $\boldsymbol{p}_{i}^{v}=\eta(\boldsymbol{z}_{i}^{v})$. Simply replace all the high-dimensional embedding $\boldsymbol{h}$ (the subscript is omitted for brevity) in Eq. (1) and Eq. (2) with its corresponding visual branch version $\boldsymbol{h}^{v}$ can obtain the the supervised contrastive loss $\mathcal{L}_{scon}^{v}$ and the self-supervised contrastive loss $\mathcal{L}_{ucon}^{v}$. The overall representation learning loss is balanced with $\lambda$, written as:
@@
For the parametric classification part, just replace $\boldsymbol{p}_{i}$ and ${\boldsymbol{p}_{i}}^{\prime}$ of Eq. (3) and Eq. (4) with $\boldsymbol{p}_{i}^{v}$ and ${\boldsymbol{p}_{i}^{v}}^{\prime}$ can obtain the cross-entropy loss $\mathcal{L}_{\text{cls-v}}^{s}$ and the self-distillation loss $\mathcal{L}_{\text{cls-v}}^{u}$. Thus, the classification loss is $\mathcal{L}_{\text{cls}}^{v}=(1-\lambda)\mathcal{L}_{\text{cls-v}}^{u}+\lambda\mathcal{L}_{\text{cls-v}}^{s}$.
@@
$\mathcal{L}_{\text{db}}^{v}=\mathcal{L}_{\text{rep}}^{v}+\mathcal{L}_{\text{cls}}^{v}\,.$ (10)
Text-branch.  Our text branch simply adopts the same training strategy as the visual branch. That is, in particular, given a text embedding $\boldsymbol{\hat{z}}_{i}^{t}$ generated by TES, we first input it into a fully connected layer to gain a learnable text embedding $\boldsymbol{\hat{z}}_{i}^{tl}$ while change its dimension. Simply replace $\boldsymbol{h}_{i}^{v}$ in the representation learning objective $\mathcal{L}_{\text{rep}}^{v}$ with $\boldsymbol{h}_{i}^{t}=g(\boldsymbol{\hat{z}}_{i}^{tl})$ and replace ${\boldsymbol{p}_{i}^{v}}$ in classification parts $\mathcal{L}_{\text{cls}}^{v}$ with $\boldsymbol{p}_{i}^{t}=\eta(\boldsymbol{\hat{z}}_{i}^{tl})$ can yield the corresponding text objectives $\mathcal{L}_{\text{rep}}^{t}$ and $\mathcal{L}_{\text{cls}}^{t}$. In other words, changing the visual conception indicator $v$ into text conception indicator $tl$ can get the corresponding objectives for the text branch. Thus, the overall objective of our text branch can be formally written as:
$\mathcal{L}_{\text{db}}^{t}=\mathcal{L}_{\text{rep}}^{t}+\mathcal{L}_{\text{cls}}^{t}\,.$ (11)
@@
Cross-modal instance consistency objective.  In order to enable the two branches to learn from each other while encouraging agreement between two different modes, we propose a cross-modal instance consistency objective (CICO), shown in Fig. 2 (c). Our CICO has the same form of mutual knowledge distillation as  [13], but we distill the instance consistency between the two branches. For each mini-batch $B$, we choose its labeled subset $B_{l}$ containing $K$ categories as anchor samples, calculate the visual and text prototypes for $K$ categories as visual anchors $\mathcal{P}_{v}$ and text anchors $\mathcal{P}_{t}$, respectively. We define the instance relationships in visual and text branch as:
$\begin{split}s_{i}^{v}=&\sigma({\boldsymbol{z}_{i}^{v}}^{\top}\mathcal{P}_{v})\\ s_{i}^{t}=&\sigma({\boldsymbol{\hat{z}}_{i}^{tl}}^{\top}\mathcal{P}_{t})\,.\end{split}$ (12)
@@
where $D_{KL}$ is the Kullback-Leibler divergence. Mutual knowledge distillation on instance relationships for two modalities makes visual and text flows exchange and benefit from each other, thus the two branches can serve as complementary discriminative aids to each other.